Boswell Company manufactures two products, Regular and Supreme. Boswell’s overhead costs consist of machining, $5000000; and assembling, $2200000. Information on the two products is:

Regular Supreme
Direct labor hours 10000 15000
Machine hours 10000 30000
Number of parts 90000 160000


Overhead applied to Regular using traditional costing using direct labor hours is

$4320000.
$1990000.
$5210000.
$2880000.
Answer #1

Total Overhead costs = $5000000 + $2200000 = $7200000
Total DIrect Labor hours = 10000+15000 = 25000

Overhead cost per direct labor hour = $7200000 / 25000 = $288

Overhead applied to Regular = $288 x 10000 = $2880000